THE WEAKLY-BOUND COMPLEX CH_{4}-H_{2}: OBSERVATION AND ANALYSIS OF INFRARED SPECTRA IN THE 350 AND 1311 cm^{−1} REGIONS

\underline{A.R.W. McKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada; D.A. ROTH, I. PAK, and G. WINNEWISSER, I. Physikalisches Institut, Universität zu Köln, Zulpicherstrasse 77, 50937 Köln, Germany.

Infrared spectra of the weakly-bound van der Waals complex CH_{4}-paraH_{2} have been observed and analyzed for the first time. Measurements were made using a long path (160-180 m) low temperature (61-92 K) absorption cell which was probed with a Bomem Fourier transform spectrometer (350 and 1311 cm^{−1}) or a tunable diode laser (1311 cm^{−1}). The partly resolved spectrum accompanying the S_{0}(0) pure rotational transition of H_{2} around 350 cm^{−1} was analyzed in terms of an approximate model to obtain the rotational and centrifugal distortion constants of the complex in its ground state. The high resolution spectrum accompanying the R(0) transition of the ν_{4} fundamental band of CH_{4} near 1311 cm^{−1} was assigned in detail and analyzed using the Coriolis model used previously for analogous spectra of methane - rare gas complexes.^{a} The effective intermolecular separation and dissocation energy of the complex in its ground state were determined to be about 4.12 Å and 28 cm^{−1}, respectively. Absorption due to CH_{4}-paraH_{2} has also been observed in the 1317 cm^{−1} region near the R(1) transition of CH_{4}, but this is not yet analyzed.

aI. Pak, D.A. Roth, M. Hepp, G. Winnewisser, D. Scouteris, B.J. Howard, and K.M.T. Yamada, Z. Naturforsch. 53a, 725 (1998).